# Interpretable Robot Control via Structured Behavior Trees and Large Language Models

### v1

@@ 1 Introduction
Figure 1: High-level overview of the proposed LLM-driven robotic control system, where a user interacts with the system through natural language, interpreted by an LLM to guide robot behavior via a modular control structure.
@@
Bringing these elements together, this paper proposes a framework that utilizes the interpretative power of LLMs alongside the structured execution of BTs to enable robust robot behavior in response to natural language instructions. The goal is to develop an end-to-end application that supports a dynamic “command-to-execution” flow, real-time decision-making, and seamless integration of new behaviors. It facilitates scalable integration of various robotic capabilities while maintaining transparency and flexibility in execution. As depicted in Fig. 2, the LLM bridges human objective and actionable robot behaviors by triggering domain-specific behavior modules through the behavior tree. The system is developed to support multiple behavior modules with effortless integration, primarily within the computer vision domain, including object tracking [13] and hand gesture recognition [14].
@@
  • A modular robotic framework that integrates LLMs with a BT to interpret natural language commands and execute defined actions,
  • An open-source implementation of the proposed system, facilitating reproducibility and further research; and
@@ 2.1 LLM-driven HRI Frameworks
Incorporating multimodality, ROSGPT_Vision [^2] https://github.com/bilel-bj/ROSGPT_Vision [19] extends ROSGPT by integrating visual language models and perception modules to ground language in visual context. However, its end-to-end prompt-based design, relying on a single vision module and fixed prompt logic, limits its adaptability to dynamic environments and tasks that require context-sensitive perception. LaMI [8] enhances multimodal understanding by integrating speech (audio), vision, text, and gestures to infer high-level human intent. However, its main limitation lies in its substantial computational overhead, stemming from its intricate multimodal fusion mechanisms.
@@ 3 Proposed Framework
The proposed system is a modular and open-source framework accessible via https://github.com/snt-arg/robot_suite, featuring an LLM interface built on top of ROSA’s [18] architecture. It is designed to be robot-agnostic and adaptable to a wide range of robotic tasks and hardware configurations. As shown in Fig. 3, the system works as an end-to-end application with components that can evolve independently, enabling plug-and-play integration of new plugins (i.e., behaviors) without complexity or the need for overall system modifications.
According to the figure, when the user issues natural language commands, they serve as the primary inputs for task interpretation and planning. These commands reflect the user’s intention to initiate specific robot actions, which in turn demand triggering the appropriate control drivers3.1) tailored for the robot. The unstructured inputs are then processed by the LLM interface module (§3.2), which employs GPT-4 backbone to translate them into structured action commands. At the core of the system, a behavior tree3.3) handles high-level decision-making procedures based on the input interpreted by the LLM, also offering task execution alongside transparent reasoning. To support various capabilities, a set of behavior modules3.4) (implemented as callable plugins) is integrated into the system. These behavior modules (e.g., person-following) are independently developed and can be triggered by the BT and operate in conjunction with hardware-specific drivers. In this regard, the LLM interface interacts with both the behavior tree (by activating relevant nodes of the tree) and the behavior modules, either by querying their status or through bidirectional communication (as in the person-following task, where it requires to specify the target and receive on-demand updates). It should be noted that although we show a sample set of executable plugins and the behavior tree structure, the system supports a diverse range of behaviors, making it extensible to exploit more robot capabilities.
@@ 3.1 Robots and Drivers
Interfacing between the high-level decision logic and the physical robots, the Drivers module plays a crucial role in executing commands generated by the system (either the LLM interface or the plugins). These drivers implement atomic, hardware-specific operations (such as take_off, land, or stand), tailored to the capabilities of different robot platforms. The driver layer abstracts and encapsulates these low-level controls, enabling straightforward integration across different robots. The current version of the framework supports both legged robots, such as the Boston Dynamics Spot®, and aerial platforms, including drones like the DJI Tello. However, it is designed to be easily extensible to additional robot platforms with minimal effort required for integration. As shown in Fig. 3, robot-specific drivers are continuous interfaces between the physical robot hardware and the rest of the framework, as they facilitate bidirectional communications by executing low-level control commands (e.g., motion and velocity) and transmitting robot status feedback (e.g., battery or connectivity).
For the supported robots $\mathbf{R}=\{r_{i}\mid i\in\mathbb{N}\}$, there exists a mapping $\mathcal{D}:\mathbf{R}\rightarrow\mathcal{D}_{R}$ such that each robot $r_{i}\in\mathbf{R}$ is associated with a specific driver $d_{r_{i}}=\mathcal{D}(r_{i})$. Accordingly, task execution on robot $r_{i}$ is delegated to the corresponding driver $d_{r_{i}}$ via an interface function $\mathcal{I}:\mathcal{C}\times\mathbf{R}$, where $\mathcal{C}=\{c_{j}\mid j\in\mathbb{N}\}$ is a set of low-level robot commands (e.g., land). Thus, the appropriate low-level action on $r_{i}$ is triggered using:
@@ 4 Experimental Results
A series of real-world experiments in diverse scenarios was conducted to evaluate the proposed system’s capabilities and potential. In this regard, we design realistic scenarios, collect our in-house dataset, and evaluate the system’s performance in various tasks. It is worth noting that the proposed system incorporates a diverse range of behaviors; however, for the experiments, we employed a limited but representative behavior tree structure along with a curated set of sample behavior modules.
@@ 4.1 Evaluation Criteria
[…]
@@
  • Success Rate: the degree to which the system achieves the intended task outcome across scenarios $\Phi 1$ to $\Phi 6$. Given the variability in LLM responses, we adopt a weighted success strategy, where each case of scenarios is evaluated over $k=10$ independent runs and across the introduced stages $\rho$. Each demanded behavior is scored using a binary metric $\eta^{\rho}_{i,j}\in\{0,1\}$, where $i$ indexes the scenario case and $j$ the iteration, with score $1$ denoting success and 0 denoting failure. Thus, success rate is calculated as:
@@
It should be noted that explicitly evaluating the performance of the underlying LLM and computer vision models is not in the primary scope of this paper. These components are treated as back-ends, and their benchmarking will be explored as a direction for future work.
@@ 4.4 Runtime Performance
In most scenarios, the majority of the time is spent querying the LLM in $\mathcal{L}_{\text{cog}}$, which can vary depending on the task complexity and the length of the generated text. Notably, $\Phi 3.5$ requires a longer time due to the amount of information that must be displayed to the user. Regarding $\mathcal{L}_{\text{exec}}$, most actions require the standard execution time of the platform. The highest execution time can be seen in $\Phi 6.1$, where the robot remains in an infinite loop while following the person. Moreover, Motion Command scenarios in $\Phi 4.\mathrm{X}$ exhibit great execution time due to multiple factors, including the magnitude of the requested motion value (e.g., distance or altitude) and hardware-specific capabilities. While the first factor directly affects how long the robot must move before completing the task, the latter is due to the robot’s locomotion variation. For instance, the legged robot is inherently slower than the drone, resulting in longer completion times.
@@ 5 Conclusions
This paper presented an open-source, modular, and extensible framework that combines large language models with behavior trees, enabling robots to interpret and execute natural language commands in a transparent and scalable manner. The core contributions of the work lie in simplifying human-robot interaction and the interpretability of robot decision-making, enabling more dynamic and adaptable robot behaviors. Real-world experiments across diverse environments showed that the system demonstrates a high success rate in both “semantic interpretation” of user instructions and “execution of corresponding robotic behaviors” stages. These experiments were structured into defined scenarios, encompassing perception-driven tasks and motion commands, to evaluate the system’s robustness and versatility under realistic conditions.
Future work will address system design aspects not covered in this paper. First, we plan to benchmark the impact of various LLM back-ends (e.g., OpenAI, SentenceTransformer, Llama, etc.) and perception-driven behavior modules on the system’s decision-making and execution quality. Additionally, extending the framework to multi-robot coordination is another step to evaluate collaborative task execution across heterogeneous robots. Finally, integrating more input modalities and supporting dynamic behavior selection through the BT are other extensions for further enhancing the adaptability of the framework.

### v2

@@ 1 Introduction
Figure 1: High-level overview of the proposed LLM-driven robotic control method, where a user interacts with the system through natural language, interpreted by an LLM to guide robot behavior via a modular control structure.
@@
Bringing these elements together, this paper proposes a novel end-to-end framework that unifies the interpretative power of LLMs alongside the structured execution of BTs to enable robust robot behavior in response to natural language instructions. Our goal, as illustrated in Fig. 1, is to present an end-to-end framework that supports a dynamic “command-to-execution” flow, real-time decision-making, and seamless integration of new behaviors. In contrast to existing state-of-the-art methods, our approach emphasizes modularity, interpretability, and scalability, by autonomously translating natural language commands into executable actions. As depicted in Fig. 2, the LLM bridges human objective and actionable robot behaviors by triggering domain-specific behavior modules through the behavior tree. The system is developed to support multiple behavior modules with effortless integration, primarily within the computer vision domain, including object tracking [13] and hand gesture recognition [14].
@@
  • A modular and model-agnostic robotic framework that integrates LLMs with customized BTs to interpret natural language commands and autonomously execute defined actions,
  • An open-source implementation of the complete framework, promoting reproducibility and further research; and
@@ 2.1 LLM-driven HRI Frameworks
Incorporating multimodality, ROSGPT_Vision [^2] https://github.com/bilel-bj/ROSGPT_Vision [19] extends ROSGPT by integrating visual language models and perception modules to ground language in a visual context. However, its end-to-end prompt-based design, relying on a single vision module and fixed prompt logic, limits its adaptability to dynamic environments and tasks that require context-sensitive perception. LaMI [8] enhances multimodal understanding by integrating speech (audio), vision, text, and gestures to infer high-level human intent. However, its main limitation lies in its substantial computational overhead, stemming from its intricate multimodal fusion mechanisms.
@@ 3 Proposed Framework
The proposed system is a modular and open-source framework publicly accessible via https://github.com/snt-arg/robot_suite, featuring an LLM interface built on top of ROSA’s [18] architecture. It is designed to be robot-agnostic and adaptable to a wide range of robotic tasks and hardware configurations. As shown in Fig. 3, the system operates as an end-to-end application with components that can evolve independently, enabling plug-and-play integration of new plugins (i.e., behaviors) without complexity or the need for overall system modifications.
According to the figure, when the user issues natural language commands, they serve as the primary inputs for task interpretation and planning. These commands reflect the user’s intention to initiate specific robot actions, which in turn demand triggering the appropriate control drivers3.1) tailored for the robot. The unstructured inputs are then processed by the LLM interface module (§3.2), which employs GPT-4 backbone to translate them into structured action commands. At the core of the system, a behavior tree3.3) handles high-level decision-making procedures based on the input interpreted by the LLM, also offering task execution alongside transparent reasoning. To support various capabilities, a set of behavior modules3.4) (implemented as callable plugins) is integrated into the system. These behavior modules (e.g., person-following) are independently developed and can be triggered by the BT, operating in conjunction with hardware-specific drivers. In this regard, the LLM interface interacts with both the behavior tree (by activating relevant nodes of the tree) and the behavior modules, either by querying their status or through bidirectional communication (as in the person-following task, where it requires specifying the target and receiving on-demand updates). It should be noted that although we present a sample set of executable plugins and the behavior tree structure, the system supports a diverse range of behaviors, making it extensible to employ additional robot capabilities.
@@ 3.1 Robots and Drivers
Interfacing between the high-level decision logic and the physical robots, the Drivers module plays a crucial role in executing commands generated by the system (either the LLM interface or the plugins). These drivers implement atomic, hardware-specific operations (such as take_off, land, or stand), tailored to the capabilities of different robot platforms. The driver layer abstracts and encapsulates these low-level controls, enabling straightforward integration across robots. The current version of the framework supports both legged robots, such as the Boston Dynamics Spot®, and aerial platforms, including drones like the DJI Tello. However, it is designed to be easily extensible to additional robot platforms with minimal effort required for integration. As shown in Fig. 3, robot-specific drivers provide the interface layer between the physical robot hardware and the rest of the framework, handling bidirectional communications by executing low-level control commands (e.g., motion and velocity) and transmitting robot status feedback (e.g., battery or connectivity).
Consequently, for the supported robots $\mathbf{R}=\{r_{i}\mid i\in\mathbb{N}\}$, there exists a mapping $\mathcal{D}:\mathbf{R}\rightarrow\mathcal{D}_{R}$ such that each robot $r_{i}\in\mathbf{R}$ is associated with a specific driver $d_{r_{i}}=\mathcal{D}(r_{i})$. Accordingly, task execution on robot $r_{i}$ is delegated to the corresponding driver $d_{r_{i}}$ via an interface function $\mathcal{I}:\mathcal{C}\times\mathbf{R}$, where $\mathcal{C}=\{c_{j}\mid j\in\mathbb{N}\}$ is a set of low-level robot commands (e.g., land). Thus, the appropriate low-level action on $r_{i}$ is triggered using:
@@ 4 Experimental Results
A series of real-world experiments in diverse scenarios was conducted to evaluate the proposed system’s capabilities and potential. In this regard, we designed realistic scenarios, collected our in-house dataset, and evaluated the system’s performance in various tasks. It is worth noting that the proposed system incorporates a diverse range of behaviors; however, for the experiments, we employed a limited but representative behavior tree structure along with a curated set of sample behavior modules.
@@ 4.1 Evaluation Criteria
A battery level below 2020% (E) is considered low, while levels above are treated as sufficient (C). Category # Platform User Instruction Environment Status Expected System Behavior & Response Unsupported ($\Phi 1$) $\Phi 1.1$ [Uncaptioned image]/[Uncaptioned image] “Jump” landed/stand/flying Not triggered, response: “I cannot perform this action.” Context-Aware Response ($\Phi 2$) $\Phi 2.1$ [Uncaptioned image] “Do a Flip” flying, C Flip maneuver executed. $\Phi 2.2$ [Uncaptioned image] “Do a Flip” landed, C Blocked due to status, response: “I cannot do it as the drone is on the ground.” $\Phi 2.3$ [Uncaptioned image] “Do a Flip” landed, E Blocked due to battery level and status, response: “I cannot do it due to low battery and robot status.” System Inquiry Response ($\Phi 3$) $\Phi 3.1$ [Uncaptioned image] “What is the battery level?” landed/flying, C Response: “The battery level is X%”. $\Phi 3.2$ [Uncaptioned image] “Can I do a flip with the drone?” flying, C Response: “Yes, since the drone is flying and the battery level is 26%”. $\Phi 3.3$ [Uncaptioned image] “Which actions can I perform with the drone?” landed/flying, C Response: “You can [a list of supported actions]”. $\Phi 3.4$ [Uncaptioned image] “What is the status of the drone?” landed/flying, C Response: “The drone is on the ground with a battery of 26%”. $\Phi 3.5$ [Uncaptioned image] “What are the common causes that I get unknown status?” landed, disconnected Response: “The common causes for the robot state being unknown could be [a list of common causes]”. Motion Command ($\Phi 4$) $\Phi 4.1$ [Uncaptioned image]/[Uncaptioned image] “Turn left/right for X sec.” stand/flying, C Motion action triggered with duration control. $\Phi 4.2$ [Uncaptioned image]/[Uncaptioned image] “Move forward/backward for X sec (with velocity Y).” stand/flying, C Motion action triggered with duration control. Plugin Switching ($\Phi 5$) $\Phi 5.1$ [Uncaptioned image]/[Uncaptioned image] “Change the control to hand gesture.” stand/flying, C Control mode switched, response: “You can now control the robot using hand gestures.” $\Phi 5.2$ [Uncaptioned image]/[Uncaptioned image] “Change the control to keyboard.” stand/flying, C Control mode switched, response: “You can now control the robot using the keyboard.” Vision-based Interaction ($\Phi 6$) $\Phi 6.1$ [Uncaptioned image] “Track the person with a phone” flying, C Person identified and approached, response: “Now tracking the person with a phone.” $\Phi 6.2$ [Uncaptioned image] “Track the person with a phone” flying, C Blocked due to identification failure; the drone waits for $5~\mathrm{sec.}$ to find the person and then halts, response: “No person with a phone detected”
@@
  • Success Rate: the degree to which the system achieves the intended task outcome across scenarios $\Phi 1$ to $\Phi 6$. Given the variability in LLM responses, we adopt a weighted success strategy, where each scenario case is evaluated over $k=10$ independent runs and across the introduced stages $\rho$. Each demanded behavior is scored using a binary metric $\eta^{\rho}_{i,j}\in\{0,1\}$, where $i$ indexes the scenario case and $j$ the iteration, with score $1$ denoting success and 0 denoting failure. Thus, success rate is calculated as:
@@
It should be noted that explicitly benchmarking the performance of underlying LLM and computer vision models is beyond the current scope of this paper. The primary objective of the proposed framework is to remain model-agnostic and to support any standard LLM or vision model that provides an accessible API for request–response communication. In this design, the LLM and perception components act as interchangeable back-end modules responsible for language understanding and environmental perception, respectively. This modular architecture enables the integration of other models without requiring modifications to the overall system structure. While the selection of a particular LLM (such as OpenAI GPT in our implementation) may influence factors such as reasoning quality or response latency, our focus in this work is on demonstrating the end-to-end functionality, scalability, and generalizability of the framework. A systematic comparison of alternative LLM back-ends and their influence on overall system performance will be explored in future work.
@@ 4.3 Success Rate Analysis
In summary, the experimental results demonstrate that the proposed framework maintains a high level of reliability and semantic consistency across heterogeneous robotic platforms and diverse interaction scenarios. Scenarios with lower $\sigma$ confirm that a clear trend is highly observable between cognition and execution success, where lower $\eta_{i}^{\text{cog}}$ values mainly correspond to higher reasoning ambiguity or context misinterpretation. This suggests that the quality of the underlying language model’s semantic understanding is the primary factor influencing the system’s end-to-end performance.
@@ 4.4 Runtime Performance
In most scenarios, the majority of the time is spent querying the LLM in $\mathcal{L}_{\text{cog}}$, which can vary depending on the task complexity and the length of the generated text. In other words, $\mathcal{L}_{\text{cog}}$ dominates the overall response time due to the time required by the LLM to generate reasonable responses. Notably, $\Phi 3.5$ requires a longer time due to the amount of information that must be displayed to the user. Regarding $\mathcal{L}_{\text{exec}}$, most actions require the standard execution time of the platform. The highest execution time can be seen in $\Phi 6.1$, where the robot remains in an infinite loop while following the person. Moreover, Motion Command scenarios in $\Phi 4.\mathrm{X}$ exhibit great execution time due to multiple factors, including the magnitude of the requested motion value (e.g., distance or altitude) and hardware-specific capabilities. While the first factor directly affects how long the robot must move before completing the task, the latter is due to the robot’s locomotion variation. For instance, the legged robot is inherently slower than the drone, resulting in longer completion times.
In general, despite the inherent variability in LLM processing, the system consistently exhibits stable timing behavior and minimal dispatch overhead, demonstrating the efficiency of the $\mathrm{ROS}$-based communication layer and the modular system architecture. Additionally, scenarios with higher cognition times are primarily associated with task complexity and context-dependent reasoning, where the LLM needs to process longer or more ambiguous instructions. This is not directly aligned with the $\mathcal{L}_{\text{exec}}$ results, as the execution time is mainly influenced by the robot’s locomotion characteristics and platform-specific speed.
@@ 5 Conclusions
This paper presented an open-source, modular, and extensible framework that combines LLMs with behavior trees, enabling robots to interpret and execute natural language commands in a transparent and scalable manner. The core contributions of this work lie in simplifying human-robot interaction and the interpretability of robot decision-making, enabling more dynamic and adaptable robot behaviors. Real-world experiments across diverse environments have shown that the system demonstrates a high success rate in both the “semantic interpretation” of user instructions and the “execution of corresponding robotic behaviors” stages. These experiments were structured into defined scenarios, encompassing perception-driven tasks and motion commands, to evaluate the system’s robustness and versatility under realistic conditions.
Future work will address system design aspects not covered in this paper. First, we plan to benchmark the impact of various LLM back-ends (e.g., OpenAI, ST, Llama) and perception-driven behavior modules on the system’s decision-making and execution quality. Additionally, extending the framework to multi-robot coordination is another step to evaluate collaborative task execution across heterogeneous robots. Finally, integrating more input modalities and supporting dynamic behavior selection through the BT are additional extensions that can further enhance the framework.